@ Consider an economy that only produces two goods: strawberries and cream. Use the table below to compute the nominal GDP, real GDP, the GDP deflator for each year, and the inflation rate. The year 2015 is the base year. SHOW ALL WORK.  

Year Price of strawberries (per pint) The number of strawberries (pints) Price of cream (per pint) The quantity of cream (pints)
2015 $3.00 100 $2.00 200
2016 $4.00 125 $2.50 400
2017 $5.00 150 $3.00 500

@ Define Gross Domestic Product and then list and define all of its components. BE SPECIFIC
Answer #1

(1)

(A) Nominal GDP (NGDP) = \sum(Current year price x Current year quantity)

2015 ($) = 3 x 100 + 2 x 200 = 300 + 400 = 700

2016 ($) = 4 x 125 + 2.5 x 400 = 500 + 1000 = 1500

2017 ($) = 5 x 150 + 3 x 500 = 750 + 1500 = 2250

(B) Real GDP (RGDP) = \sum(Base year (2015) price x Current year quantity)

2015 ($) = 3 x 100 + 2 x 200 = 300 + 400 = 700

2016 ($) = 3 x 125 + 2 x 400 = 375 + 800 = 1175

2017 ($) = 3 x 150 + 2 x 500 = 450 + 1000 = 1450

(C) GDP Deflator = (NGDP / RGDP) x 100

2015 = ($700 / $700) x 100 = 100

2016 = ($1500 / $1175) x 100 = 127.66

2017 = ($2250 / $1450) x 100 = 155.17

(D) Inflation rate = % Change in GDP Deflator

2016 = (127.66 / 100) - 1 = 1.2766 - 1 = 0.2766 = 27.66%

2017 = (155.17 / 127.66) - 1 = 1.2155 - 1 = 0.2155 = 21.55%

NOTE: As per Answering Policy, 1st question is answered.